A block is kept on a smooth inclined plane of angle of inclinaton 30 degree that moves with a constant acceleration so that the block does not slide relative to the inclined plane .Let F1 be the contact force between the block and the plane .Now the inclined plane stops and let F2 be the contact force between the two in this case .Then F1/F2 IS 
(a) 1       (b)4/3    (c)2    (d)3/2

Grade:12th pass

1 Answers

abinayaa
5 years ago
in the first case,
F1 sin = 30=ma
and F1 cos=30=mg or f1 = mg/cos30
 
and in the second case,
f2 =mg cos 30
 
therefore,
F2 = mg cos 30
F1/F2 =1/cos^2 30 = 4/3